Circular No. 6981

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

NOVA IN M31 M. Modjaz, E. Halderson, T. Shefler, J. Y. King, W. D. Li, R. R.

Treffers, and A. V. Filippenko, University of California at Berkeley, report their discovery of an apparent nova during the course of the Lick Observatory Supernova Search (cf. IAUC 6627) with the 0.8-m Katzman Automatic Imaging Telescope (KAIT). Available magnitude estimates from unfiltered CCD images: July 23.5 UT, ∼ 16.1; 27.5, ∼ 14.4; 31.5, ∼ 15.6. The object is located at α = 0^{h}42^{m}37^{s}.00, δ = +41^{o}16′56′′.7 (equinox 2000.0), which is ∼ 82′′.4 west and 47′′.7 north of the nucleus of NGC 224 (M31). A KAIT image on July 19.5 does not show a star at the position of the apparent nova (limiting mag ∼ 18.0).

SUPERNOVA 1998dh IN NGC 7541 Magnitude estimates, visual unless otherwise noted: July 20.98 UT,

16.8 (S. Massaro, Palermo, Italy, CCD; prediscovery image); 26.48, 14.4 (N. Biver, Oahu, HI); 27.49, 14.2 (Biver); 30.99, 14.60 (D. Rodriguez, Madrid, Spain, CCD + V filter); Aug. 2.0, 14.3 (K. Sárneczky, Ágasvár, Hungary).

S/1997 U 1

Extension to the ephemeris on IAUC 6870:

1998 TT α_{2000} δ_{2000} Δ r_{U} ϵ Δα′′ Δδ′′
July 26 20^{h}54.^{m}96 −18^{o}01.^{′}8 18.822 0.046 171.^{o}7 85− 226+
Aug. 5 20 53.40 −18 07.8 18.816 0.047 178.2 45− 264+
15 20 51.84 −18 13.7 18.839 0.047 168.3  3− 298+
25 20 50.36 −18 19.2 18.892 0.047 158.3 38+ 329+
Sept. 4 20 49.03 −18 24.2 18.973 0.048 148.3 80+ 354+
14 20 47.88 −18 28.3 19.080 0.048 138.4 121+ 375+
24 20 46.98 −18 31.5 19.209 0.049 128.4 160+ 391+
Oct. 4 20 46.37 −18 33.7 19.356 0.049 118.4 197+ 402+
14 20 46.08 −18 34.6 19.518 0.050 108.5 232+ 408+
24 20 46.12 −18 34.2 19.689 0.050 98.5 264+ 410+
Nov. 3 20 46.50 −18 32.6 19.865 0.050 88.7 293+ 408+
13 20 47.22 −18 29.7 20.041 0.051 78.8 318+ 402+
23 20 48.27 −18 25.6 20.211 0.051 69.0 339+ 393+
Dec. 3 20 49.61 −18 20.3 20.371 0.051 59.2 357+ 380+
13 20 51.23 −18 14.0 20.515 0.051 49.4 370+ 364+

1998 August 3  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Daniel W. E. Green